# AccompGen: Hierarchical Autoregressive Vocal Accompaniment Generation with Dual-Rate Codec Tokenization

### 10 Apr 2026 v1

@@ Abstract
We present AccompGen, a system that generates instrumental music audio to accompany input vocals. Given isolated singing voice, AccompGen produces a coherent instrumental accompaniment that can be directly mixed with the input to create complete music. We propose three key innovations over prior work: (1) a dual-rate codec tokenization scheme using HuBERT semantic tokens at 50 Hz for vocals and EnCodec acoustic tokens at 75 Hz for instrumentals, enabling time-aligned yet rate-independent modeling; (2) a three-stage hierarchical autoregressive architecture (semantic $\rightarrow$ coarse acoustic $\rightarrow$ fine acoustic) with interleaved multi-codebook prediction and classifier-free guidance; and (3) modern Transformer design choices including QK-norm, GEGLU activations, RMSNorm, and T5-style relative position bias for improved training stability and sequence generalization. Experiments on MUSDB18 demonstrate that AccompGen achieves a Fréchet Audio Distance (FAD) of 2.02 on isolated vocal inputs, outperforming retrieval baselines and matching prior state-of-the-art systems with fewer parameters.
@@ 1 Introduction
In this work, we present AccompGen, which advances vocal accompaniment generation through three contributions:
@@
Fig. 1: Overview of AccompGen. Left: Data preprocessing pipeline. Source separation extracts aligned (vocal, instrumental) pairs from music mixtures. Vocals are encoded to HuBERT semantic tokens (50 Hz); instrumentals to EnCodec acoustic tokens (75 Hz). Right: Three-stage hierarchical AR generation. Stage 1 predicts instrumental semantic tokens from vocal semantics. Stage 2 predicts coarse acoustic codes (4 codebooks, interleaved) conditioned on both vocal and instrumental semantics. Stage 3 predicts fine acoustic codes (4 codebooks) from coarse codes. EnCodec decodes the full 8-codebook sequence to the instrumental waveform.
@@ 3.2 Dual-Rate Codec Tokenization
3.3 Three-Stage Hierarchical AR Model
We decompose generation into three stages, each modeled by an independent decoder-only causal Transformer:
Stage 1 (Semantic): $P(\mathbf{s}^{i}\mid\mathbf{s}^{v})$ — predicts instrumental semantic tokens from vocal semantic tokens. Both operate at 50 Hz with vocabulary size 500.
Stage 2 (Coarse): $P(\mathbf{a}^{c}\mid\mathbf{s}^{v},\mathbf{s}^{i})$ — predicts 4-codebook coarse acoustic codes at 75 Hz, conditioned on vocal and instrumental semantics. Multi-codebook tokens are interleaved: $q^{1}_{1},q^{2}_{1},q^{3}_{1},q^{4}_{1},q^{1}_{2},q^{2}_{2},\ldots]$, with per-quantizer offset embeddings to disambiguate codebook identity.
Stage 3 (Fine): $P(\mathbf{a}^{f}\mid\mathbf{a}^{c})$ — predicts 4-codebook fine acoustic codes from coarse codes, refining audio fidelity.
The overall generation proceeds as:
@@ 3.4 Transformer Architecture
Each stage uses a shared architecture with the following design:
@@
QK-Norm. We apply RMSNorm independently to queries and keys before computing attention scores [7], preventing attention logit growth during training:
@@
where RN denotes RMSNorm and $\mathbf{B}$ is the relative position bias.
@@
with inner dimension $d_{\text{ff}}=\lfloor\frac{2}{3}\cdot d_{\text{model}}\cdot\text{mult}\rfloor$ following the two-thirds convention.
T5 Relative Position Bias. Instead of absolute positional embeddings, we use learned relative position biases with logarithmic bucketing (32 buckets) [14], shared across all layers. This enables generalization to longer sequences at inference.
@@ 3.5 Training
Classifier-free guidance (CFG). During training, we randomly drop conditioning inputs with probability $p_{\text{drop}}\!=\!0.1$ per sample. At inference, we interpolate between conditional and unconditional logits:
$\hat{\ell}_{t}=(1+\lambda)\cdot\ell_{t}^{\text{cond}}-\lambda\cdot\ell_{t}^{\text{uncond}},$ (5)
where $\lambda\!=\!3.0$ is the guidance scale.
Loss function. Each stage is trained with cross-entropy loss. For multi-codebook stages, the loss is averaged over quantizers:
@@
Training configuration. We use AdamW [13] ($\beta_{1}\!=\!0.9$, $\beta_{2}\!=\!0.98$), cosine LR schedule with 4000-step warmup, learning rate $3\times 10^{-4}$, FP16 mixed precision, and gradient clipping at norm 1.0. Model dimensions: $d_{\text{model}}\!=\!512$, $n_{\text{heads}}\!=\!8$, $n_{\text{layers}}\!=\!12$, $d_{\text{ff\_mult}}\!=\!4.0$.
@@ 3.6 Data Preprocessing
We apply source separation (MDXNet [12]) to a large corpus of music audio to extract aligned (vocal, instrumental) pairs. Following SingSong [8], we filter clips where the instrumental is silent (<25<-25 dB RMS) or the vocal exceeds the instrumental by more than 5 dB, biasing the system toward always producing audible instrumentals.
@@
Stage Batch Clip (s) Steps Grad Acc. Seq Len
Semantic 180 10 100K 1 1000
Coarse 36 5 200K 2 1500
Fine 120 3 150K 1 900
@@ 4.2 Quantitative Results
AccompGen-Base (Ours) 2.02 2.30 0.33
@@ 5 Conclusion
We presented AccompGen, a three-stage hierarchical autoregressive system for vocal accompaniment generation. By combining dual-rate HuBERT/EnCodec tokenization, interleaved multi-codebook AR modeling with CFG, and modern Transformer components, AccompGen achieves strong results on MUSDB18 while using fewer parameters than comparable systems. Future work includes scaling to higher sample rates, supporting multi-source generation (e.g., separate drum, bass, piano tracks), and conditioning on additional attributes such as genre and style.

### 13 Apr 2026 v2

# HAFM: Hierarchical Autoregressive Foundation Model for Music Accompaniment Generation

@@ Abstract
We present HAFM, a system that generates instrumental music audio to accompany input vocals. Given isolated singing voice, HAFM produces a coherent instrumental accompaniment that can be directly mixed with the input to create complete music. We propose three key innovations over prior work: (1) a dual-rate codec tokenization scheme using HuBERT semantic tokens at 50 Hz for vocals and EnCodec acoustic tokens at 75 Hz for instrumentals, enabling time-aligned yet rate-independent modeling; (2) a three-stage hierarchical autoregressive architecture (semantic $\rightarrow$ coarse acoustic $\rightarrow$ fine acoustic) with interleaved multi-codebook prediction and classifier-free guidance; and (3) modern Transformer design choices including QK-norm, GEGLU activations, RMSNorm, and T5-style relative position bias for improved training stability and sequence generalization. Experiments on MUSDB18 demonstrate that HAFM achieves a Fréchet Audio Distance (FAD) of 2.08 on isolated vocal inputs, outperforming retrieval baselines and matching prior state-of-the-art systems with fewer parameters. The source code is available at https://github.com/HackerHyper/HAFM.
@@ 1 Introduction
In this work, we present HAFM, which advances vocal accompaniment generation through three contributions:
@@
Fig. 1: Data preprocessing pipeline. Source separation (MDXNet) extracts aligned vocal and instrumental pairs from music mixtures. Vocals are augmented with Gaussian noise ($\sigma\!=\!0.01$) and encoded via HuBERT-Large (layer 9) with $k$-means (k=500k\!=\!500) into semantic tokens at 50 Hz. Instrumentals are encoded into both semantic tokens (HuBERT, 50 Hz) and acoustic tokens via EnCodec (8 codebooks at 75 Hz), split into coarse (CB 1–4) and fine (CB 5–8) groups.
@@ 3.2 Dual-Rate Codec Tokenization
3.3 Three-Stage Hierarchical Autoregressive Model
We decompose generation into three stages, each modeled by an independent decoder-only causal Transformer. All stages share the same architecture hyperparameters but differ in their input/output token specifications. During training, each stage receives conditioning tokens and teacher-forced target tokens concatenated as input, with classifier-free guidance (CFG) dropout applied to conditioning. The training procedure for all three stages is illustrated in Fig. 2.
Stage 1 (Semantic): $P(\mathbf{s}^{i}\mid\mathbf{s}^{v})$ — predicts instrumental semantic tokens from vocal semantic tokens (Fig. 2(a)). Both operate at 50 Hz with vocabulary size 500. The vocal semantic tokens serve as the sole conditioning input. This stage captures the high-level musical correspondence between vocals and instrumentals.
Stage 2 (Coarse): $P(\mathbf{a}^{c}\mid\mathbf{s}^{v},\mathbf{s}^{i})$ — predicts 4-codebook coarse acoustic codes at 75 Hz, conditioned on both vocal and instrumental semantics (Fig. 2(b)). Multi-codebook tokens are interleaved into a flat sequence: for Q=4Q\!=\!4 codebooks and $T$ time frames, the prediction target is $a^{1}_{1},a^{2}_{1},a^{3}_{1},a^{4}_{1},a^{1}_{2},a^{2}_{2},\ldots,a^{Q}_{T}]$, yielding a sequence of length $Q\times T$. Per-quantizer offset embeddings disambiguate codebook identity within the interleaved sequence. The conditioning tokens from each source are embedded separately and concatenated with distinct segment IDs ($N_{\text{seg}}\!=\!3$: one per conditioning source plus one for prediction).
Stage 3 (Fine): $P(\mathbf{a}^{f}\mid\mathbf{a}^{c})$ — predicts 4-codebook fine acoustic codes conditioned on coarse codes (Fig. 2(c)), refining audio fidelity. The coarse conditioning uses Q=4Q\!=\!4 codebook embeddings that are summed across quantizers at each time step, projecting the multi-codebook input into a single embedding per frame.
The overall inference pipeline (Fig. 3) proceeds as:
@@ 3.3 Three-Stage Hierarchical Autoregressive Model
(a) Stage 1: Semantic prediction. Conditioning: vocal semantic tokens $\mathbf{s}^{v}$. Target: instrumental semantic tokens $\mathbf{s}^{i}$ (50 Hz, V=500V\!=\!500). Loss: cross-entropy over 1 codebook.
(b) Stage 2: Coarse acoustic prediction. Conditioning: $\mathbf{s}^{v}$ and $\mathbf{s}^{i}$. Target: interleaved coarse codes $\mathbf{a}^{c}$ (4 CB $\times$ 75 Hz). Loss: cross-entropy averaged over 4 quantizers.
(c) Stage 3: Fine acoustic prediction. Conditioning: coarse codes $\mathbf{a}^{c}$. Target: fine codes $\mathbf{a}^{f}$ (4 CB $\times$ 75 Hz). Loss: cross-entropy averaged over 4 quantizers.
Fig. 2: Training procedure for the three-stage hierarchical autoregressive model. Each stage is trained independently with teacher forcing and CFG dropout ($p_{\text{drop}}\!=\!0.1$). Conditioning and target tokens are embedded with separate token and segment embeddings, concatenated, and fed through a causal Transformer. Per-quantizer linear heads produce output logits.
Fig. 3: Inference pipeline of HAFM. Given a vocal waveform, HuBERT encodes it into semantic tokens $\mathbf{s}^{v}$. Stage 1 autoregressively generates instrumental semantic tokens $\mathbf{s}^{i}$ with CFG ($\lambda\!=\!3.0$). Stage 2 produces coarse acoustic codes $\mathbf{a}^{c}$ conditioned on both $\mathbf{s}^{v}$ and $\mathbf{s}^{i}$. Stage 3 refines $\mathbf{a}^{c}$ into fine codes $\mathbf{a}^{f}$. The 8-codebook codes are decoded by EnCodec to produce the instrumental waveform, which is mixed with the original vocal to create complete music.
@@ 3.4 Transformer Architecture
Each stage uses a shared decoder-only causal Transformer architecture (Fig. 4) with L=12L\!=\!12 layers, $d_{\text{model}}\!=\!512$, $n_{\text{heads}}\!=\!8$ ($d_{k}\!=\!64$), and $d_{\text{ff\_mult}}\!=\!4.0$. We incorporate several modern design choices for improved training stability and sequence generalization.
Input Embeddings. Each conditioning source has its own token embedding $\mathbf{E}_{\text{cond}}\in\mathbb{R}^{(V+1)\times d}$, where the extra index serves as a mask token for CFG dropout. For multi-codebook conditioning (Stage 3), per-quantizer embeddings are summed to produce a single vector per frame. Prediction tokens use per-quantizer embeddings $\mathbf{E}_{\text{pred}}^{q}\in\mathbb{R}^{(V_{p}+1)\times d}$ ($q\!=\!1,\ldots,Q$), where the extra index is the BOS token. A learned quantizer-offset embedding $\mathbf{E}_{\text{qoff}}\in\mathbb{R}^{Q\times d}$ is added to interleaved prediction positions to signal codebook identity. A segment embedding $\mathbf{E}_{\text{seg}}\in\mathbb{R}^{N_{\text{seg}}\times d}$ distinguishes conditioning sources from prediction tokens. All embeddings are initialized from $\mathcal{N}(0,0.02)$.
@@
A final RMSNorm is applied to the Transformer output before the prediction heads.
QK-Norm. We apply RMSNorm independently to queries and keys (per-head, on dimension $d_{k}\!=\!64$) before computing attention scores [7], preventing attention logit growth during training:
@@
where RN denotes RMSNorm and $\mathbf{B}$ is the relative position bias. Queries, keys, and values are computed jointly via a single fused projection $\mathbf{W}_{QKV}\in\mathbb{R}^{d\times 3d}$ without bias.
@@
with inner dimension $d_{\text{ff}}=\lfloor\frac{2}{3}\cdot d_{\text{model}}\cdot\text{mult}\rfloor=1365$, following the two-thirds convention from LLaMA [3]. The gate and value branches share the input projection ($\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{d\times d_{\text{ff}}}$), and $\mathbf{W}_{3}\in\mathbb{R}^{d_{\text{ff}}\times d}$ projects back to model dimension.
T5 Relative Position Bias. Instead of absolute positional embeddings, we use learned relative position biases with logarithmic bucketing [14]. The bias is computed via 32 buckets with causal-only configuration: the first 16 buckets encode exact offsets (01515), while the remaining 16 cover log-spaced distances up to a maximum distance derived from the sequence length. The bias is shared across all layers, computed once per forward pass, and broadcast across the batch dimension.
Per-Quantizer Output Heads. For single-codebook stages, a single linear head $\mathbf{W}_{\text{out}}\in\mathbb{R}^{d\times V}$ maps the Transformer output to logits. For multi-codebook stages, $Q$ independent linear heads $\{\mathbf{W}_{\text{out}}^{q}\}_{q=1}^{Q}$ produce per-quantizer logits. During training, each interleaved position is routed to its corresponding head via modular indexing ($q=t\bmod Q$). During inference, the active head is selected based on the current generation step.
Fig. 4: Transformer block architecture used in all three stages. Each block consists of pre-norm RMSNorm, multi-head attention with QK-Norm and T5 relative position bias, residual connections, and a GEGLU feed-forward network. Configuration: $d_{\text{model}}\!=\!512$, $n_{\text{heads}}\!=\!8$, L=12L\!=\!12 layers, $d_{\text{ff\_mult}}\!=\!4.0$.
@@ 3.5 Training
Each stage is trained independently with teacher forcing. As shown in Fig. 2, conditioning tokens and right-shifted target tokens are embedded with per-segment and per-quantizer embeddings, concatenated along the sequence dimension, and fed through the causal Transformer. A BOS token (the extra vocabulary index $V_{p}$) is prepended to the target sequence for autoregressive training, so the model predicts position $t$ from all positions <t<t.
Classifier-free guidance (CFG). During training, we randomly drop conditioning inputs with probability $p_{\text{drop}}\!=\!0.1$ per sample, replacing all conditioning token embeddings with zeros for the dropped samples. This trains the model to generate both conditionally and unconditionally. At inference, we perform two forward passes per step and interpolate:
$\hat{\ell}_{t}=\ell_{t}^{\text{uncond}}+\lambda\cdot(\ell_{t}^{\text{cond}}-\ell_{t}^{\text{uncond}}),$ (5)
where $\lambda\!=\!3.0$ is the guidance scale. This doubles the inference cost but significantly improves generation quality.
Loss function. Each stage is trained with cross-entropy loss. For single-codebook Stage 1, the loss is computed directly over the full prediction sequence. For multi-codebook Stages 2 and 3, per-quantizer linear heads produce independent logits and the loss is averaged over $Q$ quantizers:
@@
Inference sampling. At each autoregressive step, the model applies temperature scaling ($\tau\!=\!0.9$), top-$k$ filtering (k=250k\!=\!250), and multinomial sampling. For the cross-rate Stage 2, the number of acoustic frames is computed as $T_{a}=T_{v}\times 75/50$ to maintain temporal alignment between 50 Hz semantic and 75 Hz acoustic rates.
Training configuration. We use AdamW [13] ($\beta_{1}\!=\!0.9$, $\beta_{2}\!=\!0.98$, weight decay $0.01$), cosine LR schedule with 4000-step linear warmup (peak $3\times 10^{-4}$, minimum $10^{-6}$), FP16 mixed precision with gradient scaling, and gradient clipping at norm 1.0. Training uses DDP across 7 A40 GPUs with NCCL backend. Stage-specific configurations are listed in Table 1.
@@ 3.6 Data Preprocessing
As illustrated in Fig. 1, we apply source separation (MDXNet [12]) to a large corpus of music audio to extract aligned (vocal, instrumental) pairs. Following SingSong [8], we filter clips where the instrumental is silent (<25<-25 dB RMS) or the vocal exceeds the instrumental by more than 5 dB, biasing the system toward always producing audible instrumentals.
@@
Stage Batch Clip (s) Steps Grad Acc. Seq Len $N_{\text{seg}}$
Semantic 210 10 100K 1 1000 2
Coarse 42 5 200K 2 1500 3
Fine 140 3 150K 1 900 2
@@ 4.2 Quantitative Results
HAFM-Base (Ours) 2.08 2.30 0.33
@@ 5 Conclusion
We presented HAFM, a three-stage hierarchical autoregressive system for vocal accompaniment generation. By combining dual-rate HuBERT/EnCodec tokenization, interleaved multi-codebook AR modeling with CFG, and modern Transformer components, HAFM achieves strong results on MUSDB18 while using fewer parameters than comparable systems. Future work includes scaling to higher sample rates, supporting multi-source generation (e.g., separate drum, bass, piano tracks), and conditioning on additional attributes such as genre and style.